OpenStudy (samsan9):

what is the equation of the hyperbola with vertices at (-4,0) and foci at (-5,0) and (5,0)

OpenStudy (samsan9):

i believe that is this

OpenStudy (samsan9):

\[\frac{ x^2 }{ 16 }-\frac{ y^2 }{ 9 }=1 \]

OpenStudy (samsan9):

but i want to make sure